# Transfer learning from first-principles calculations to experiments with chemistry-informed domain transformation

### v1

@@ Abstract
Simulation-to-Real (Sim2Real) transfer learning, the machine learning technique that efficiently solves a real-world task by leveraging knowledge from computational data, has received increasing attention in materials science as a promising solution to the scarcity of experimental data. We proposed an efficient transfer learning scheme from first-principles calculations to experiments based on the chemistry-informed domain transformation, that integrates the heterogeneous source and target domains by harnessing the underlying physics and chemistry. The proposed method maps the computational data from the simulation space (source domain) into the space of experimental data (target domain). During this process, these qualitatively different domains are efficiently bridged by prior knowledge of chemistry, the statistical ensemble and the relationship between source and target quantities.
@@ 1 Introduction
In practice, the data-driven approach for materials often fails because of the lack of real data. Experimental data in materials science are scarce and non-scalable due to several reasons: (1) the high cost and time required for synthesis and measurement; (2) the disparate modality depending on measurement methods; and (3) the exploration bias, where the data collected is biased towards known or easily accessible regions of the material space, making ML tasks extrapolate beyond existing data. Although there are attempts to obtain large datasets by running high-throughput experiments [6, 7, 8, 9, 10], combinatorial synthesis [11, 12, 13, 14, 15], and laboratory automation [16, 17, 18], the available materials and conditions are limited, and the data sizes are still insufficient, typically of the order of $O(100)$, below the requirement of common DL methods.
On the other hand, computational data generated by first-principles calculations such as density functional theory (DFT) have been increasingly used in the context of ML. Compared to real experiments, numerical computations are scalable and easily automated, providing a large and abundant dataset. Such computational datasets obtained by high-throughput DFT calculations are available in a wide variety of materials [19, 20, 21, 22, 23, 24, 25]. However, computational data also has limitations. Computational methods rely on several approximations and assumptions, and thus systematic errors must appear in the results. The impact of such errors on predictions in reality cannot be known without comparison with experiments. Although it is sometimes possible to reproduce reality by performing high-fidelity calculations, such calculations are often expensive and significantly degrade its data throughput. Therefore, computational and experimental data are complementary to each other in which the former is lower cost but lower fidelity and the other is vice versa, motivating the introduction of transfer learning. Such a transfer scheme, called the Sim2Real (Simulation-to-Real) transfer, has been successfully introduced in some areas [26, 27, 28, 29].
Although it is reasonable to hypothesize that the Sim2Real transfer is also effective for material discovery, this is challenging due to the fundamental differences between first-principles calculations and real experiments. The most fatal issue is the difference in the scales: a first-principles calculation provides a microscopic description of a single (often simple) structure, which is represented by a set of atoms; an experimental measurement captures a macroscopic profile of a composite of various structures distributed near thermal equilibrium, which is usually difficult to control or even characterize.
@@ 2.3 Physics/Chemistry-Informed Machine Learning
Similarly, chemistry-informed machine learning leverages chemical laws, which are mainly based on thermodynamics and statistical mechanics. Bradford et al. [44] developed a prediction model for ionic conductivity in polymer electrolytes based on a chemistry-informed neural network. Their architecture incorporates the Arrhenius equation in the final layers, which describes the temperature-activated processes, leading to a significant improvement. Ballard et al. [45] performed machine learning modeling for polymerization processes with a chemistry-informed neural network. They incorporate the knowledge of kinetic models into the loss function, resulting in an improved model outperforming conventional ones.
@@ 3.1 Overview
Figure 2: Schematics of the chemistry-informed domain transformation. The rectangle box represent the process, rounded rectangle represent the data and its domain, the doubled-rounded rectangle represents the chemical information (CI), respectively.
@@ 3.2 Domain transformation procedure
We next show a detailed procedure of the chemistry-informed domain transformation as shown in Figure 2 Let the domain be $\mathbb{D}^{r}\equiv({\mathcal{X}}^{r}\times{\mathcal{Y}}^{r},P^{r}_{X,Y})$ with the feature space ${\mathcal{X}}^{r}$, the output space ${\mathcal{Y}}^{r}$, and the joint probability distribution $P^{r}_{X,Y}$ where $r={\mathrm{S}}$ or ${\mathrm{T}}$ represent the source and target domain, respectively. Then the domain transformation aims to construct transformation rules for the source domain as follows:
@@
These procedures result in $\mathbb{D}^{\mathrm{S}}\to\mathbb{D}^{{\mathrm{S}}^{\prime}}\approx\mathbb{D}^{{\mathrm{T}}}$, reducing the problem to a homogeneous domain shift. In particular, from the construction of the domain transformation, if the output space $\mathcal{Y}$ is perfectly aligned, it is expected that the relationship known as covariate shift [31] will hold between the transformed computational data and the experimental data. That is, we can expect that the relation $P^{{\mathrm{S}}^{\prime}}_{Y|X}\approx P^{\mathrm{T}}_{Y|X}$ approximately holds. Covariate shift is a typical model for the discrepancy between different domains, and various correction methods have been proposed in machine learning comunity [46, 47]. By utilizing these methods, it is possible to achieve effective learning of prediction models under the covariate shift.
@@ 3.2.1 Ensemble averaging:
We first transform the feature space as ${\mathcal{X}}^{\mathrm{S}}\times{\mathcal{Y}}^{\mathrm{S}}\to{\mathcal{X}}^{\mathrm{T}}\times{\mathcal{Y}}^{\mathrm{S}}$. The objective of this step is to obtain projected source data in the target feature space, $\tilde{\mathcal{D}}^{\mathrm{S}}=\{(x^{{\mathrm{S}}^{\prime}},\tilde{y}^{{\mathrm{S}}})\}$ where $(x^{{\mathrm{S}}^{\prime}},\tilde{y}^{{\mathrm{S}}})\in{\mathcal{X}}^{\mathrm{T}}\times{\mathcal{Y}}^{\mathrm{S}}$. Here, $x^{{\mathrm{S}}^{\prime}}\in{\mathcal{X}}^{\mathrm{T}}$ is an auxiliary variable with the same feature space as $x^{\mathrm{T}}$. $\tilde{y}^{{\mathrm{S}}}\in{\mathcal{Y}}^{\mathrm{S}}$ is the ensemble average of $y^{{\mathrm{S}}}$ associated with $x^{{\mathrm{S}}^{\prime}}$:
@@
where $k_{\mathrm{B}}$ is the Boltzmann constant, $E(x^{{\mathrm{S}}})$ is the energy of $x^{{\mathrm{S}}}$. This is known as the canonical ensemble. As an advanced example, let us consider the case of varying the chemical composition under isothermal conditions. In this case, $x^{{\mathrm{S}}^{\prime}}$ is a chemical composition and $\theta=T$ is the fixed temperature, leading:
@@ 3.2.3 Function estimation:
In this work, we assume that there exists a conversion function $F:{\mathcal{Y}}^{\mathrm{S}}\to{\mathcal{Y}}^{\mathrm{T}}$ from a source quantity to a target one. This $F$ describes some relation between $\tilde{y}^{{\mathrm{S}}}$ and $y^{{\mathrm{T}}}$. It may be either a black box function, an analytic function, or a non-parametric function, depending on the problem. The goal of this step is to estimate $F$ by using $\mathcal{D}^{{\mathrm{S}}\cap{\mathrm{T}}}$.
@@ 4.1 Catalyst activity prediction for RWGS
On the other hand, from the computational perspective, the activation energy cannot be computed directly because it requires the knowledge of the entire reaction path, which is computationally unfeasible. Instead, first-principles calculations commonly provide the adsorption energy, the binding energy between an adsorbate and a catalyst surface, which is relatively easy to obtain from structural relaxation [50]. Although this quantity is closely related to the activation energy, they are qualitatively different because the adsorption process is only a part of the total reaction process.
@@ 11                        $f_{M}^{\mathrm{occ}}\leftarrow f_{M}^{\mathrm{occ}}+1$;
Figure 6: Linear regression of activation energy by adsorption energy of the sources selected with S2 strategy.
Applying the Algorithm 2 to $\{\mathcal{D}^{{\mathrm{S}}(M)\cap{\mathrm{T}}}\}_{M}$, Figure 5 displays the adsorbates whose occurrence frequency exceeds 50%. Theirs linear regression lines are shown in Figure 6, presenting different lines depending on the sources. We consider these to be candidate sources and expect any one of them to give an adequate conversion. The selected sources from the strategy S2 are summarized in Table. 1. Detailed results and numerical values of the function estimation are given in C.1.
@@ 4.3.4 Conversion of physical quantity:
We convert $\tilde{\bm{y}}_{i}^{{\mathrm{S}}}$ to ${y}^{{\mathrm{S}}^{\prime}}$ with the estimated function $\hat{F}$, resulting in source data on the target space ${\mathcal{D}}^{{\mathrm{S}}^{\prime}}=\{({x}^{{\mathrm{S}}^{\prime}},{y}^{{\mathrm{S}}^{\prime}})\}$, with the transformed source domain, $\mathbb{D}^{{\mathrm{S}}^{\prime}}=({\mathcal{X}}^{\mathrm{T}}\times{\mathcal{Y}}^{\mathrm{T}},P_{X,Y}^{{\mathrm{S}}^{\prime}})$. Figures 7 compare the label distribution of the transformed source data ($P(Y^{{\mathrm{S}}^{\prime}})$) and target data ($P(Y^{\mathrm{T}})$) for Wang2023 and the dummy data, respectively. In all Adsorbates, the transformed source distribution is closer in shape to the target distribution than the raw OC20 distributions in Figures 7. Furthermore, to quantitatively evaluate the effectiveness of our domain transformation, we calculated the discrepancy between distributions of the source and target domain before and after the transformation using the Wasserstein distance [55]. The Wasserstein distance can be interpreted as a measure of the cost required to transform one probability distribution into another and a smaller Wasserstein distance suggests that the two distributions are closer to each other. Table 2 shows the Wasserstein distances between the source (C2O1) distribution and the target distribution, suggesting that the distributions are closer to each other by domain transformation.
@@ 4.3.5 Homogeneous transfer learning:
Finally, transfer learning pert performs the homogeneous domain adaptation. For input of the model, the composition is converted into the chemical descriptors generated by the XenonPy code [56]. We employ 224 descriptors that is relevant to our problem out of 290 descriptor, shown in Appendix B. These descriptors are standardized using all compositions in OC20 dataset.
@@ 4.5.2 Demonstration with real data:
Next, let us evaluate performances for real experimental data from Wang2023 by using the pretrained model with OC20 (C2O1) dataset. We also find positive transfer in this case from Figure 10, indicating approximately 1/10 of the error reduction by knowledge transfer. Compared to Figure 9, the error at the start is higher, but still be lower than any value of the loss of scratch model; Thus, the transferred knowledge provides near-optimal parameters on the target domain. Similar to the previous section, the FT and the TL show almost same performance.
@@ 4.5.3 Comparison of source datasets:
At the end of this section, we will investigate source data dependencies. Since transfer method does not affect the model performance as shown in above, TL is used in this section.
We evaluated the performances of transferred model for different source datasets in OC20 selected with the strategy S1 or S2, summarized in Table 3. Figure 11 shows the model performance for some of the sources included in the table. This result indicates that the final test errors depends on the source selection, however, transferred model performs enough better than the scratch model regardless of the used source. Although C2O1 from S1 strategy performs relatively better in terms of final accuracy, H3N1 from S2 strategy reaches similar accuracy.
@@ 4.6 Discussions
Our demonstrations in both a dummy dataset and a real dataset indicate that the proposed method successfully transfers the knowledge from source computational datasets beyond the domain gap, enabling highly efficient learning on the target real-world task. Pretraining with transformed source data, using only a few (1̃0) target data for domain transformation, shows a significant improvement in prediction accuracy with reducing test losses one order of magnitude.
Moreover, the results show that the transferred knowledge gives nearly optimal parameters in target model, suggesting a potential to reduce a number of experiments required to develop prediction models for novel catalysts.
@@ 5 Conclusions
Through a verification with dummy target data and a demonstration with the experimental data, we confirmed that the proposed method shows positive transfer of increasing accuracy and data efficiency compared to the model trained only with target data. The test losses of transferred model was approximately one order of magnitude smaller than that of full scratch model. Moreover, the source model has already been optimized in the target space, despite only using less than 10 target data during the domain transformation.
@@ A.3 Dummy data
First, a prediction model $\hat{h}$ trained by Wang2023 is prepared. Because only $\mathcal{D}^{{\mathrm{T}}}$ is not enough to train with a neural network-based model, we here employ $\hat{h}$ as a gradient boosting decision tree model implemented in LightGBM code [60]. To train the model, compositions are converted into the XenonPy descriptor (See Appendix. B). The essential parameters are listed in Table 5.
@@ Appendix B Composition featurizer
The XenonPy compositional featurizer is a tool within the XenonPy library that converts raw chemical composition data into informative numerical descriptors suitable for machine learning [56]. It can calculate 290 compositional features for a given chemical composition. This calculation uses the information of the 58 elemental properties. Using weights of each element in a composition, the 290 compositional descriptors are obtained from the five calculations, weighted sum, weighted average, weighted variance, max-pooling, and min-pooling, on the 58 elemental properties ($58\times 5=290$). For more detail, see the code repository.
In this work, since all $f$-block elements are excluded from our datasets in advance, we ignore num_f_unfilled and num_f_valence out of the 58 elemental properties. Moreover, we omit the weighted sum calculation as it is identical to the weighted average for a periodic system. Consequently, we use $56\times 4=224$ descriptors.

### v2

@@ Abstract
Simulation-to-Real (Sim2Real) transfer learning, the machine learning technique that efficiently solves a real-world task by leveraging knowledge from computational data, has received increasing attention in materials science as a promising solution to the scarcity of experimental data. We proposed an efficient transfer learning scheme from first-principles calculations to experiments based on the chemistry-informed domain transformation that bridges the heterogeneous source and target domains by harnessing the underlying physics and chemistry. The proposed method maps the computational data from the simulation space (source domain) into the space of experimental data (target domain). During this process, these qualitatively different domains are efficiently integrated by a couple of prior knowledge of chemistry, (1) the statistical ensemble, and (2) the relationship between source and target quantities.
@@ 1 Introduction
In practice, the data-driven approach for materials often fails because of the lack of real data. Experimental data in materials science are scarce and non-scalable due to several reasons: (1) the high cost and time required for synthesis and measurement; (2) the disparate modality depending on measurement methods; and (3) the exploration bias, where the data collected is biased towards known or easily accessible regions of the material space, making ML tasks extrapolate beyond existing data. Although there are attempts to obtain large datasets by running high-throughput experiments [6, 7, 8, 9, 10], combinatorial synthesis [11, 12, 13, 14, 15], and laboratory automation [16, 17, 18], the available materials and conditions are limited, and the data sizes are still insufficient, typically of the order of $O(100)$, below the requirements of common DL methods.
On the other hand, computational data generated by first-principles calculations such as density functional theory (DFT) have been increasingly used in the context of ML. Compared to real experiments, numerical computations are scalable and easily automated, providing a large and abundant dataset. Such computational datasets obtained by high-throughput DFT calculations are available in a wide variety of materials [19, 20, 21, 22, 23, 24, 25]. However, computational data also has limitations. Computer simulations rely on several approximations and assumptions, and thus systematic errors must appear. The impact of such errors cannot be known without comparison with experiments. Although it is sometimes possible to reproduce a part of reality by performing high-fidelity calculations, such calculations are often expensive and significantly degrade their data throughput. Therefore, computational and experimental data are complementary to each other in which the former is lower cost but lower fidelity and the other is vice versa, motivating the introduction of transfer learning. In fact, a transfer learning framework, called the Sim2Real (Simulation-to-Real) transfer, has already been successfully introduced in some areas [26, 27, 28, 29].
Although it is reasonable to hypothesize that the Sim2Real transfer is also effective for material discovery, this is still challenging due to the fundamental differences between first-principles calculations and real experiments. The most fatal issue is the difference in the scales: a first-principles calculation provides a microscopic description of a single (often simple) structure, which is represented by a set of atoms; an experimental measurement captures a macroscopic profile of a composite of various structures distributed near thermal equilibrium, which is usually difficult to control or even characterize.
@@ 2.3 Physics/Chemistry-Informed Machine Learning
Similarly, chemistry-informed machine learning leverages chemical laws, which are mainly based on thermodynamics and statistical mechanics. Bradford et al. [44] developed a prediction model for ionic conductivity in polymer electrolytes based on a chemistry-informed neural network. Their architecture incorporates the Arrhenius equation in the final layers, which describes the temperature-activated processes, leading to a significant improvement. Ballard et al. [45] performed machine learning modeling for polymerization processes with a chemistry-informed neural network. They incorporate the knowledge of kinetic models into the loss function, resulting in an improved model that outperforms conventional ones.
@@ 3.1 Overview
Figure 2: Schematics of the chemistry-informed domain transformation. The rectangle box represent the process, rounded rectangle represent the data and its domain, the red doubled-rounded rectangle represents the chemical information (CI), respectively.
@@ 3.2 Domain transformation procedure
We next show a detailed procedure of the chemistry-informed domain transformation as shown in Figure 2. Let the domain be $\mathbb{D}^{r}\equiv({\mathcal{X}}^{r}\times{\mathcal{Y}}^{r},P^{r}_{X,Y})$ with the feature space ${\mathcal{X}}^{r}$, the output space ${\mathcal{Y}}^{r}$, and the joint probability distribution $P^{r}_{X,Y}$ where $r={\mathrm{S}}$ or ${\mathrm{T}}$ represent the source and target domain, respectively. Then the domain transformation aims to construct transformation rules for the source domain as follows:
@@
These procedures result in $\mathbb{D}^{\mathrm{S}}\to\mathbb{D}^{{\mathrm{S}}^{\prime}}\approx\mathbb{D}^{{\mathrm{T}}}$, reducing the problem to a homogeneous domain shift. In particular, from the construction of the domain transformation, if the output space $\mathcal{Y}$ is perfectly aligned, it is expected that the relationship known as covariate shift [31] will hold between the transformed computational data and the experimental data. That is, we can expect that the relation $P^{{\mathrm{S}}^{\prime}}_{Y|X}\approx P^{\mathrm{T}}_{Y|X}$ approximately holds. Covariate shift is a typical model for the discrepancy between different domains, and various correction methods have been proposed in the machine learning community [46, 47]. By utilizing these methods, it is possible to achieve effective learning of prediction models under the covariate shift.
@@ 3.2.1 Ensemble averaging:
We first transform the feature space as ${\mathcal{X}}^{\mathrm{S}}\times{\mathcal{Y}}^{\mathrm{S}}\to{\mathcal{X}}^{\mathrm{T}}\times{\mathcal{Y}}^{\mathrm{S}}$. The objective of this step is to obtain projected source data in the target feature space, $\tilde{\mathcal{D}}^{\mathrm{S}}=\{(x^{{\mathrm{S}}^{\prime}},\tilde{y}^{{\mathrm{S}}})\}$ where $(x^{{\mathrm{S}}^{\prime}},\tilde{y}^{{\mathrm{S}}})\in{\mathcal{X}}^{\mathrm{T}}\times{\mathcal{Y}}^{\mathrm{S}}$. Here, $x^{{\mathrm{S}}^{\prime}}\in{\mathcal{X}}^{\mathrm{T}}$ is an auxiliary variable with the same feature space as $x^{\mathrm{T}}$. $\tilde{y}^{{\mathrm{S}}}\in{\mathcal{Y}}^{\mathrm{S}}$ is the ensemble average of $y^{{\mathrm{S}}}$ associated with $x^{{\mathrm{S}}^{\prime}}$, represented by:
@@
where $k_{\mathrm{B}}$ is the Boltzmann constant, $E(x^{{\mathrm{S}}})$ is the energy of $x^{{\mathrm{S}}}$. This is known as the canonical ensemble.
As an advanced example, let us consider the case of varying the chemical composition under isothermal conditions. In this case, $x^{{\mathrm{S}}^{\prime}}$ is a chemical composition and $\theta=T$ is the fixed temperature, leading:
@@ 3.2.3 Function estimation:
In this work, we assume that there exists a conversion function $F:{\mathcal{Y}}^{\mathrm{S}}\to{\mathcal{Y}}^{\mathrm{T}}$ from a source quantity to a target one. This $F$ describes some relation between $\tilde{y}^{{\mathrm{S}}}$ and $y^{{\mathrm{T}}}$. It may be either a black box function, an analytic function, or a non-parametric function, depending on the problem. The goal of this step is to estimate optimal $\hat{F}$ by using $\mathcal{D}^{{\mathrm{S}}\cap{\mathrm{T}}}$.
@@ 4.1 Catalyst activity prediction for RWGS
On the other hand, from the computational perspective, the activation energy cannot be computed directly because it requires the knowledge of the entire reaction path, which is computationally unfeasible. Instead, first-principles calculations commonly provide the adsorption energy, the binding energy between an adsorbate and a catalyst surface, which is relatively easy to obtain from structural relaxation [50]. Although this quantity is closely related to the activation energy, they are qualitatively different because the adsorption process is only a part of the total reaction processes.
@@ 11                        $f_{M}^{\mathrm{occ}}\leftarrow f_{M}^{\mathrm{occ}}+1$;
Figure 6: Linear regression of activation energy by adsorption energy of the sources selected with the S2 strategy.
Applying the Algorithm 2 to $\{\mathcal{D}^{{\mathrm{S}}(M)\cap{\mathrm{T}}}\}_{M}$, Figure 5 displays the adsorbates whose occurrence frequency exceeds 50%. Their linear regression lines are shown in Figure 6, presenting different lines depending on the sources. We consider these to be candidate sources and expect any one of them to give an adequate conversion. The selected sources from the strategy S2 are summarized in Table 1. Detailed results and numerical values of the function estimation are given in C.1.
@@ 4.3.4 Conversion of physical quantity:
We convert $\tilde{\bm{y}}_{i}^{{\mathrm{S}}}$ to ${y}^{{\mathrm{S}}^{\prime}}$ with the estimated function $\hat{F}$, resulting in source data on the target space ${\mathcal{D}}^{{\mathrm{S}}^{\prime}}=\{({x}^{{\mathrm{S}}^{\prime}},{y}^{{\mathrm{S}}^{\prime}})\}$, with the transformed source domain, $\mathbb{D}^{{\mathrm{S}}^{\prime}}=({\mathcal{X}}^{\mathrm{T}}\times{\mathcal{Y}}^{\mathrm{T}},P_{X,Y}^{{\mathrm{S}}^{\prime}})$. Figures 7 compare the label distribution of the transformed source data ($P(Y^{{\mathrm{S}}^{\prime}})$) and target data ($P(Y^{\mathrm{T}})$) for Wang2023 and the dummy data, respectively. In all adsorbates, the transformed source distribution is closer in shape to the target distribution than the raw OC20 distributions in Figures 7. Furthermore, to quantitatively evaluate the effectiveness of our domain transformation, we calculated the discrepancy between distributions of the source and target domain before and after the transformation using the Wasserstein distance [55]. The Wasserstein distance can be interpreted as a measure of the cost required to transform one probability distribution into another, and a smaller Wasserstein distance suggests that the two distributions are closer to each other. Table 2 shows the Wasserstein distances between the source (C2O1) distribution and the target distribution, suggesting that the distributions are closer to each other by domain transformation.
@@ 4.3.5 Homogeneous transfer learning:
Finally, transfer learning pert performs the homogeneous domain adaptation. For input of the model, the composition is converted into the chemical descriptors generated by the XenonPy code [56]. We employ 224 descriptors that is relevant to our problem out of 290 descriptor, shown in B. These descriptors are standardized using all compositions in OC20 dataset.
@@ 4.5.2 Demonstration with real data:
Next, let us evaluate performances for real experimental data from Wang2023 by using the pretrained model with OC20 (C2O1) dataset. We also find positive transfer in this case from Figure 10, indicating approximately 1/10 of the error reduction by knowledge transfer. Compared to Figure 9, the error at the start is higher, but still lower than any value of the loss of the scratch model; Thus, the transferred knowledge provides near-optimal parameters on the target domain. Similar to the previous section, the FT and the TL show almost the same performance.
@@ 4.5.3 Comparison of source datasets:
At the end of this section, we will investigate source data dependencies. Since the transfer method does not affect the model performance as shown above, TL is used in this section.
We evaluated the performances of the transferred model for different source datasets in OC20 selected with the strategy S1 or S2, summarized in Table 3. Figure 11 shows the model performance for some of the sources included in the table. This result indicates that the final test errors depend on the source selection; however, the transferred model performs better than the scratch model regardless of the used source. Although C2O1 from the S1 strategy performs relatively better in terms of final accuracy, H3N1 from the S2 strategy reaches similar accuracy.
@@ 4.6 Discussions
Our demonstrations in both a dummy dataset and a real dataset indicate that the proposed method successfully transfers the knowledge from source computational datasets beyond the domain gap, enabling highly efficient learning on the target real-world task. Pretraining with transformed source data, using only a few (about 10) target data for domain transformation, shows a significant improvement in prediction accuracy with reduced test losses by one order of magnitude.
Moreover, the results show that the transferred knowledge gives nearly optimal parameters in the target model, suggesting a potential to reduce the number of experiments required to develop prediction models for novel catalysts.
@@ 5 Conclusions
Through a verification with dummy target data and a demonstration with the experimental data, we confirmed that the proposed method shows positive transfer of increasing accuracy and data efficiency compared to the model trained only with target data. The test losses of transferred model was approximately one order of magnitude smaller than that of full scratch model. Moreover, the source model has already been optimized in the target space, despite only using about 10 target data during the domain transformation.
@@ A.3 Dummy data
First, a prediction model $\hat{h}$ trained by Wang2023 is prepared. Because only $\mathcal{D}^{{\mathrm{T}}}$ is not enough to train with a neural network-based model, we here employ $\hat{h}$ as a gradient boosting decision tree model implemented in LightGBM code [60]. To train the model, compositions are converted into the XenonPy descriptor (See B). The essential parameters are listed in Table 5.
@@ Appendix B Composition featurizer
The XenonPy compositional featurizer is a tool within the XenonPy library that converts raw chemical composition data into informative numerical descriptors suitable for machine learning [56]. It can calculate 290 compositional features for a given chemical composition. This calculation uses the information of the 58 elemental properties. The compositional descriptors are obtained from the five calculations, weighted-sum, weighted-average, weighted-variance, max-pooling, and min-pooling, using the 58 elemental properties and the composition ratio of each element ($58\times 5=290$). For more details, see the code documentation (https://xenonpy.readthedocs.io/en/latest/features.html).
In this work, since all $f$-block elements are excluded from our datasets in advance, we ignore num_f_unfilled and num_f_valence out of the 58 elemental properties. Moreover, we omit the weighted-sum calculation as it is identical to the weighted-average for a periodic system. Consequently, we use $56\times 4=224$ descriptors.